OpenStudy (anonymous):

MEDAL ALERT!!!!!!!!!!!! If f(x) = 7x - 1, what does f(12) represent? The value if x when (7x - 1) = 12 The value of (7x - 1) when x = 12 The value of 12 f(x) when x = 12 The value of f(x) when (7x - 1) = 12

OpenStudy (anonymous):

what you are doing in f(12)is plugging 12 in for x

OpenStudy (anonymous):

he's right

OpenStudy (anonymous):

i kind just gave the answer away agian didnt i?

OpenStudy (anonymous):

what do you think the answer is?

OpenStudy (anonymous):

d maybe?

OpenStudy (anonymous):

when 7x - 1 = 12 7x = 12 + 1 x = 13 / 7

OpenStudy (anonymous):

C then?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

no becasue what you are saying is like y=y. however true it does not answer the question

OpenStudy (anonymous):

the value of ( 7x - 1 ) when x = 12 7 ( 12 ) - 1 84 -1 83

OpenStudy (anonymous):

so d?

OpenStudy (anonymous):

wit im sorry that was wrong what c is really saying is 12y=y. so that is wrong

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so d then

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

or a

OpenStudy (anonymous):

because of the function?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

but zack says no

OpenStudy (anonymous):

w3hy

OpenStudy (anonymous):

i mean B

OpenStudy (anonymous):

BWOHOOOOO

OpenStudy (anonymous):

sorry because f(12) = translating 12 to x

OpenStudy (anonymous):

o ok oooooh I SEEE !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

i wish i can medal u both for your help

OpenStudy (anonymous):

who wants the medal whoever answers k first

OpenStudy (anonymous):

Just say k

OpenStudy (anonymous):

for the medal

OpenStudy (anonymous):

thanks man